OpenStudy (anonymous):

Let u = <-3, -5>, v = <-3, 1>. Find u + v. <-2, -8> <-8, -2> <0, -6> <-6, -4> @Shay17 @douglaswinslowcooper

OpenStudy (anonymous):

D. last choice is the only one that is correct !

OpenStudy (anonymous):

-3 + -3 and -5 + 1

OpenStudy (anonymous):

how did you find it

OpenStudy (anonymous):

how did you find the answer

OpenStudy (anonymous):

I just told you

OpenStudy (anonymous):

-3+-3=-6 and -5+1=-4 so the answer is (-6,-4)

OpenStudy (anonymous):

It asked you to find u+v so add x+x and y+y

OpenStudy (anonymous):

oh i see thanks could you help me in this one please Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v. what does the u stand for and the v ? thanks

OpenStudy (anonymous):

They are just the variable of the coordinate that lets you know that (-5,1) is u and (7,-4) is v

OpenStudy (anonymous):

And if you want help post it as a new question

OpenStudy (anonymous):

Yes. if u=(a,b) and v=(c,d) then u+v = (a+c, b+d)

OpenStudy (anonymous):

so you mean i have to add -5+-4

OpenStudy (anonymous):

Don't start getting confused here.. where are you getting -4 from ?

OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v. wait -5+7 1+-4